# Derivative-enhanced Deep Operator Network

### v1

@@ Abstract
Deep operator networks (DeepONets), a class of neural operators that learn mappings between function spaces, have recently been developed as surrogate models for parametric partial differential equations (PDEs). In this work we propose a derivative-enhanced deep operator network (DE-DeepONet), which leverages the derivative information to enhance the prediction accuracy, and provide a more accurate approximation of the derivatives, especially when the training data are limited. DE-DeepONet incorporates dimension reduction of input into DeepONet and includes two types of derivative labels in the loss function for training, that is, the directional derivatives of the output function with respect to the input function and the gradient of the output function with respect to the physical domain variables. We test DE-DeepONet on three different equations with increasing complexity to demonstrate its effectiveness compared to the vanilla DeepONet.
Keywords: Neural operators, Operator learning, Derivative learning, DeepONet, dimension reduction, Adjoint method
@@ 1 Introduction
Using neural networks to approximate the maps between functions spaces governed by parametric PDEs can be very beneficial in solving many-query problems, typically arising from Bayesian inference, optimization under uncertainty, and Bayesian optimal experimental design. Indeed, once pre-trained on a dataset, neural networks are extremely fast to evaluate given unseen inputs, compared to traditional numerical methods like the finite element method. Recently many architectures are proposed to enhance the learning capacity, such as deep operator network (DeepONet) [11], physics-informed DeepONet [20], variational energy based DeepONet (V-DeepONet) [6], Fourier neural operator (FNO) [7], graph-based neural operator (GNO) [8], convolutional neural operator (CNO) [17]. Though these work demonstrate to be successful in approximating the output function, they do not necessarily provide accurate approximation of the derivative of the output with respect to the input, which typically needs to be computed many times in many problems in science and engineering such as PDE-constrained optimization problems for control, inference, and experimental design [13, 5].
In this paper, we propose a novel method to enhance the performance of DeepONet through derivative-based dimension reduction and the incorporation of derivative information in the training. For the dimension reduction, we examine two projection bases: Karhunen–Loève Expansion (KLE) basis and active subspace method (ASM) basis. We incorporate derivative information by introducing additional losses, enabling the model to learn the operator more effectively when training data are limited.
Our work is related to the derivative-informed projected neural network (DIPNet) [16] and derivative-informed neural operator (DINO) [15] and presents an extension to enhance the performance of the DeepONet. The notable differences include: Unlike the encoder-decoder architecture employed in their work, we use the DeepONet architecture in constructing the neural network. Additionally, while the output in their papers are considered as the observations of the solution at a finite number of points, our approach can approximate the full solution field. Another difference is that we also incorporate the spatial derivatives in the training.
Our main contributions can be summarized as follows: (1) We propose a novel method that significantly improves DeepONet’s approximation accuracy for the output function, its directional derivative with respect to the input function, as well as its spatial derivatives, especially when the training samples are limited. (2) We provide the details on the computation of derivative labels of the solution of PDEs in a general form as well as the derivative-based dimension reduction to largely reduce the computational cost. (3) We demonstrate the effectiveness (enhanced approximation accuracy and computation cost) of our method through experiments conducted on three distinct PDEs with increasing complexities.
The rest of the paper is presented as follows: In Section 2 we present preliminaries of the problem setup, high fidelity approximation, and DeepONet, followed by Section 3 on the detailed presentation of DE-DeepONet and demonstration experiments in Section 4. Discussion of the conclusion and limitation is presented in the last section.
@@ 2 Preliminaries
This section presents the problem setup, high-fidelity approximation using finite element, and the vanilla DeepONet.
@@ 2.2 High Fidelity Approximation
2.3 Vanilla DeepONet
We briefly review the DeepONet architecture [11] with a focus on learning the solution operator of the PDE (1). To predict the evaluation of solution function $u$ at any point $x\in\Omega\cup\partial\Omega$ for any given input function $m$, [11] design a network architecture that comprises two separate neural networks: a trunk net $t(\cdot\;;\theta_{t})$, which takes the coordinate values of the point $x$ at which we want to evaluate the solution function as inputs, and a branch net $b(\cdot\;;\theta_{b})$, which takes the vector $\bm{m}$ encoding the parameter function $m$ as inputs. In [11], the vector $\bm{m}$ is the function evaluations at a finite set of fixed points $\{x^{(j)}\}_{j=1}^{N_{h}^{\text{in}}}\subseteq\Omega\cup\partial\Omega$, that is, $\bm{m}=(m(x^{(1)}),\cdots,m(x^{(N_{h}^{\text{in}})}))^{T}$, which corresponds to coefficient vector in the finite element approximation with Lagrange basis at the same nodes. If the solution function is scalar-valued, then both neural networks output vectors of the same dimensions. The prediction is obtained by taking the standard inner product between these two vectors and (optionally) adding a real-valued bias parameter, i.e,
@@ 3 DE-DeepONet
The vanilla DeepONet uses the input parameter and output solution pairs as the labels for the model training. The approximation for the derivative of the solution with repsect to the parameter (and the coordinate) are not necessarily accurate. However, in many downstream tasks such as Bayesian inference and experimental design, the derivatives are required. To better leverage our knowledge of the underlying PDE, we consider incorporating two additional derivative labels, that is, the Fréchet derivative $du(m;\cdot)$ and spatial derivative $\nabla_{x}u$, for the supervised training of the DeepONet, which we call derivative-enhanced DeepONet (DE-DeepONet). Let $\theta$ denote the trainable parameters in the DE-DeepONet. We propose to minimize the loss function
$\begin{split}L(\theta)&=\lambda_{1}\mathbb{E}_{m\sim\nu}\|u(m)-\hat{u}(m)\|_{L^{2}(\Omega)}^{2}\\ &+\lambda_{2}\mathbb{E}_{m\sim\nu}\|du(m;\cdot)-d\hat{u}(m;\cdot)\|_{\text{HS}}^{2}\\ &+\lambda_{3}\mathbb{E}_{m\sim\nu}\|\nabla_{x}u(m)-\nabla_{x}\hat{u}(m)\|_{L^{2}(\Omega)}^{2},\end{split}$ (3)
where the $L^{2}(\Omega)$ norm of a square integrable function $f$ is defined as $\|f\|_{L^{2}(\Omega)}^{2}=(\int_{\Omega}\|f(x)\|_{2}^{2}\;\mathrm{d}x)^{1/2}$, the Hilbert–Schmidt norm of an operator $T:H\to H$ that acts on a Hilbert space $H$ is defined as $\|T\|_{\text{HS}}^{2}=\sum_{i\in I}\|Te_{i}\|_{H}^{2}$, where $\{e_{i}\}_{i\in I}$ is an orthonormal basis of $H$. Here, $\lambda_{i}>0,i=1,2,3$ are hyperparameters that balance different loss terms.
@@
We next formally introduces our DE-DeepONet, which uses dimension reduction for the input and incorporates two types of derivatives labels, that is, the directional derivative of output with respect to the input of the branch net and the derivative of output with respect to the input of the trunk net, as additional soft constraints into the loss function.
@@ 3.1 Model Architecture
where $\theta=(\theta_{b},\theta_{t},\theta_{\text{bias}})$ are the parameters to be learned. The branch net $b(\cdot;\theta_{b})$ and trunk net $t(\cdot;\theta_{t})$ can be chosen as MLP, ResNet, etc. Note that the branch net takes a small vector of the projected parameter as input.
@@ 3.2 Loss Function
$\begin{split}L(\theta)&=\frac{\lambda_{1}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)}),u(m^{(i)})(x^{(j)}))\}_{j=1}^{N_{x}})\\ &+\frac{\lambda_{2}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\nabla_{\bm{m}}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})\Psi^{\text{in}},\Phi^{\text{out}}(x^{(j)})(\nabla_{\bm{m}}\bm{u}(\bm{m}^{(i)})\Psi^{\text{in}})\}_{j=1}^{N_{x}})\\ &+\frac{\lambda_{3}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\nabla_{x}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)}),\nabla_{x}u(m^{(i)})(x^{(j)}))\}_{j=1}^{N_{x}}),\end{split}$ (5)
@@
    are obtained by first computing the Gateaux derivatives $du(m^{(i)};\psi_{1}^{\text{in}}),\ldots,du(m^{(i)};\psi_{r}^{\text{in}})$ and then evaluating them at $x^{(j)}$. See Section 3.4 for details about the computation of the Gateaux derivative $du(m;\psi)$.
  • The third term is for learning the spatial gradient of the output function $u$ at point $x^{(j)}$. The gradient outputs $\nabla_{x}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})$ are also computed via automatic differentiation, with the partial derivative being taken with respect to the input of the trunk net. The gradient labels $\nabla_{x}u(m^{(i)})(x^{(j)})$ are computed using the gradient of the finite element basis.
  • We initialize the loss weights $\lambda_{i}=1$ for $i=1,2,3$ and choose a loss balancing algorithm called the self-adaptive learning rate annealing algorithm [19] to update them at a certain frequency. This ensures that the gradient of each loss term have similar magnitudes, thereby enabling the neural network to learn all these labels simultaneously.
@@
[…]
@@ 3.3 Dimension Reduction
Throughout the paper, we assume that the parameter functions $m$ are independently drawn from some Gaussian random field [1]. In particular, we consider the case where the covariance function is the Whittle-Matern covariance function, that is, $m\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(\bar{m},\mathcal{C})$, where $\bar{m}$ is the (deterministic) mean function and $\mathcal{C}=(\delta I-\gamma\Delta)^{-2}$ ($I$ is the identity and $\Delta$ is the Laplacian) is an operator such that the square root of its inverse, $\mathcal{C}^{-\frac{1}{2}}$, maps random function $(m-\bar{m})$ to Gaussian white noise with unit variance [9]. The parameters $\delta,\gamma\in\mathbb{R}^{+}$ jointly control the marginal variance and correlation length.
We consider two well-known linear projection bases for dimension reduction of the parameter function.
@@
When solved using the finite element method, Equation 6 is equivalent to the following linear system (See Section A.1 for the derivation.)
@@
The active subspace method is a gradient-based dimension reduction method that looks for directions in the input space contributing most significantly to the output variability [3]. In contrast to the KLE basis, the ASM basis is more computationally expensive. However, since the ASM basis captures sensitivity information in the input-output map rather than solely the variability of the input space, it typically achieves higher accuracy in predicting the output than KLE basis. We consider the case where the output is a multidimensional vector [21], representing the nodal values of the output function $u$. The ASM basis $\psi$ satisfies the differential equation
@@
Here, $du(m;\psi)$ is the Gateaux derivative of $u$ at $m\in V_{h}^{\text{in}}$ in the direction of $\psi\in V_{h}^{\text{in}}$, defined as $\lim_{\varepsilon\to 0}(u(m+\varepsilon\psi)-u(m))/\varepsilon$, and $d^{*}u(m;\cdot)$ is the adjoint of the operator $du(m;\cdot)$. When solved using finite element method,  Equation 8 is equivalent to the following linear system (See Section A.2 for the derivation.)

### v2

@@ Abstract
The deep operator networks (DeepONet), a class of neural operators that learn mappings between function spaces, have recently been developed as surrogate models for parametric partial differential equations (PDEs). In this work we propose a derivative-enhanced deep operator network (DE-DeepONet), which leverages derivative information to enhance the solution prediction accuracy and provides a more accurate approximation of solution-to-parameter derivatives, especially when training data are limited. DE-DeepONet explicitly incorporates linear dimension reduction of high dimensional parameter input into DeepONet to reduce training cost and adds derivative loss in the loss function to reduce the number of required parameter-solution pairs. We further demonstrate that the use of derivative loss can be extended to enhance other neural operators, such as the Fourier neural operator (FNO). Numerical experiments validate the effectiveness of our approach.
@@ 1 Introduction
Using neural networks to approximate the maps between functions spaces governed by parametric PDEs can be very beneficial in solving many-query problems, typically arising from Bayesian inference, optimization under uncertainty, and Bayesian optimal experimental design. Indeed, once pre-trained on a dataset, neural networks are extremely fast to evaluate given unseen inputs, compared to traditional numerical methods like the finite element method. Recently various neural operators are proposed to enhance the learning capacity, with two prominent examples deep operator network (DeepONet) [1] and Fourier neural operator (FNO) [2], which are shown to be inclusive of each other in their more general settings  [3, 4], see also their variants and other related neural operator architectures in [5, 6, 7, 8, 9, 10]. Though these work demonstrate to be successful in approximating the output function, they do not necessarily provide accurate approximation of the derivative of the output with respect to the input, which are often needed for many downstream tasks such as PDE-constrained optimization problems for control, inference, and experimental design [11, 12, 13, 14].
In this paper, we propose to enhance the performance of DeepONet through derivative-based dimension reduction for the function input inspired by [15, 16, 17, 18] and the incorporation of derivative information in the training to learn both the output and its derivative with respect to the input inspired by [19, 20]. These two derivative-enhanced approaches can significantly improve DeepONet’s approximation accuracy for the output function and its directional derivative with respect to the input function, especially when the training samples are limited. We provide details on the computation of derivative labels of the solution of PDEs in a general form as well as the derivative-based dimension reduction to largely reduce the computational cost. We demonstrate the effectiveness of our proposed method (DE-DeepONet) compared to three other neural operators, including DeepONet, FNO, and derivative-informed neural operator (DINO) [20], in terms of both test errors and computational cost. In addition, we apply derivative learning to train the FNO and also compare its performance with other methods. The code for data generation, model training and inference, as well as configurations to reproduce the results in this paper can be found at https://github.com/qy849/DE-DeepONet.
@@ 2 Preliminaries
This section presents the problem setup, high-fidelity approximation using finite element for finite dimensional discretization, and the DeepONet architecture in learning the solution operator.
@@ 2.2 High fidelity approximation
2.3 DeepONet
We briefly review the DeepONet architecture [1] with a focus on learning the solution operator of the PDE in Equation 1. To predict the evaluation of solution function $u$ at any point $x\in\Omega\cup\partial\Omega$ for any given input function $m$, [1] design a network architecture that comprises two separate neural networks: a trunk net $t(\cdot\;;\theta_{t})$, which takes the coordinate values of the point $x$ at which we want to evaluate the solution function as inputs, and a branch net $b(\cdot\;;\theta_{b})$, which takes the vector $\bm{m}$ encoding the parameter function $m$ as inputs. In [1], the vector $\bm{m}$ is the function evaluations at a finite set of fixed points $\{x^{(j)}\}_{j=1}^{N_{h}^{\text{in}}}\subseteq\Omega\cup\partial\Omega$, that is, $\bm{m}=(m(x^{(1)}),\cdots,m(x^{(N_{h}^{\text{in}})}))^{T}$, which corresponds to coefficient vector in the finite element approximation with Lagrange basis at the same nodes. If the solution function is scalar-valued, then both neural networks output vectors of the same dimensions. The prediction is obtained by taking the standard inner product between these two vectors and (optionally) adding a real-valued bias parameter, i.e,
@@ 3 DE-DeepONet
The DeepONet uses the input parameter and output solution pairs as the labels for the model training. The approximation for the derivative of the solution with respect to the parameter (and the coordinate) are not necessarily accurate. However, in many downstream tasks such as Bayesian inference and experimental design, the derivatives are required. We consider incorporating the the Fréchet derivative $du(m;\cdot)$ for the supervised training of the DeepONet, which we call derivative-enhanced DeepONet (DE-DeepONet). By doing this we hope the optimization process can improve the neural network’s ability to predict the derivative of the output function with respect to the input function. Let $\theta$ denote the trainable parameters in the DE-DeepONet. We propose to minimize the loss function
$\begin{split}L(\theta)&=\lambda_{1}\mathbb{E}_{m\sim\nu}\|u(m)-\hat{u}(m)\|_{L^{2}(\Omega)}^{2}+\lambda_{2}\mathbb{E}_{m\sim\nu}\|du(m;\cdot)-d\hat{u}(m;\cdot)\|_{\text{HS}}^{2},\end{split}$ (3)
where the $L^{2}(\Omega)$ norm of a square integrable function $f$ is defined as $\|f\|_{L^{2}(\Omega)}^{2}=(\int_{\Omega}\|f(x)\|_{2}^{2}\;\mathrm{d}x)^{1/2}$, the Hilbert–Schmidt norm of an operator $T:H\to H$ that acts on a Hilbert space $H$ is defined as $\|T\|_{\text{HS}}^{2}=\sum_{i\in I}\|Te_{i}\|_{H}^{2}$, where $\{e_{i}\}_{i\in I}$ is an orthonormal basis of $H$. Here, $\lambda_{1},\lambda_{2}>0$ are hyperparameters that balance different loss terms.
@@
We next formally introduces our DE-DeepONet, which uses dimension reduction for the input of the branch net and incorporates its output-input directional derivative labels as additional soft constraints into the loss function.
@@ 3.1 Model architecture
where $\theta=(\theta_{b},\theta_{t},\theta_{\text{bias}})$ are the parameters to be learned. The branch net $b(\cdot;\theta_{b})$ and trunk net $t(\cdot;\theta_{t})$ can be chosen as an MLP, ResNet, etc. Note that the branch net takes a small vector of the projected parameter as input. We also apply the Fourier feature embeddings [5], defined as $\gamma(x)=[\cos(Bx),\sin(Bx)]$, to the trunk net, where each entry in $B\in\mathbb{R}^{m\times d}$ is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ and $m\in\mathbb{N}^{+},\sigma\in\mathbb{R}^{+}$ are hyper-parameters.
@@ 3.2 Loss function
$\begin{split}L(\theta)&=\frac{\lambda_{1}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)}),u(m^{(i)})(x^{(j)}))\}_{j=1}^{N_{x}})\\ &+\frac{\lambda_{2}}{N_{\mathcal{D}}}\sum_{i=1}^{N_{\mathcal{D}}}\text{err}(\{(\nabla_{\bm{m}}\hat{u}(\bm{m}^{(i)};\theta)(x^{(j)})\Psi^{\text{in}},\Phi^{\text{out}}(x^{(j)})(\nabla_{\bm{m}}\bm{u}(\bm{m}^{(i)})\Psi^{\text{in}})\}_{j=1}^{N_{x}}),\end{split}$ (5)
@@
    are obtained by first computing the Gateaux derivatives $du(m^{(i)};\psi_{1}^{\text{in}}),\ldots,du(m^{(i)};\psi_{r}^{\text{in}})$ and then evaluating them at $x^{(j)}$. See Section B.3 for details about the computation of the Gâteaux derivative $du(m;\psi)$.
  • We initialize the loss weights $\lambda_{1}=\lambda_{2}=1$ and choose a loss balancing algorithm called the self-adaptive learning rate annealing algorithm [23] to update them at a certain frequency. This ensures that the gradient of each loss term have similar magnitudes, thereby enabling the neural network to learn all these labels simultaneously.
@@
For the training, the computational cost of the second term of the loss function (and its gradients) largely depends on the number of points used in each iteration. To reduce the computational and especially memory cost, we can use a subset of points, $N_{x}^{\text{batch}}=\alpha N_{x}$, where $\alpha$ is small number between 00 and $1$ (e.g., $\alpha=0.1$ in our experiments), in a batch of functions, though their locations could vary among different batches in one epoch. We find that this approach has little impact on the prediction accuracy of the model when $N_{x}$ is large enough.
@@ 3.3 Dimension reduction
Throughout the paper, we assume that the parameter functions $m$ are independently drawn from some Gaussian random field [24]. In particular, we consider the case where the covariance function is the Whittle-Matérn covariance function, that is, $m\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(\bar{m},\mathcal{C})$, where $\bar{m}$ is the (deterministic) mean function and $\mathcal{C}=(\delta I-\gamma\Delta)^{-2}$ ($I$ is the identity and $\Delta$ is the Laplacian) is an operator such that the square root of its inverse, $\mathcal{C}^{-\frac{1}{2}}$, maps random function $(m-\bar{m})$ to Gaussian white noise with unit variance [25]. The parameters $\delta,\gamma\in\mathbb{R}^{+}$ jointly control the marginal variance and correlation length.
We consider two linear projection bases for dimension reduction of the parameter function.
@@
When solved using the finite element method, Equation 6 is equivalent to the following linear system (See Section A.2 for the derivation)
@@
The active subspace method is a gradient-based dimension reduction method that looks for directions in the input space contributing most significantly to the output variability [27]. In contrast to the KLE basis, the ASM basis is more computationally expensive. However, since the ASM basis captures sensitivity information in the input-output map rather than solely the variability of the input space, it typically achieves higher accuracy in predicting the output than KLE basis. We consider the case where the output is a multidimensional vector [22], representing the nodal values of the output function $u$. The ASM basis $\psi_{i}$, $i=1,\dots,r$, are the eigenfunctions corresponding to the $r$ largest eigenvalues of the generalized eigenvalue problem
@@
Here, $du(m;\psi)$ is the Gâteaux derivative of $u$ at $m\in V_{h}^{\text{in}}$ in the direction of $\psi\in V_{h}^{\text{in}}$, defined as $\lim_{\varepsilon\to 0}(u(m+\varepsilon\psi)-u(m))/\varepsilon$, and $d^{*}u(m;\cdot)$ is the adjoint of the operator $du(m;\cdot)$. When solved using finite element method,  Equation 8 is equivalent to the following linear system (See Section A.3 for the derivation)